OpenStudy (ashleyy_smith):

Medal for correct answer (: Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 5 liters of natural oil for every 4 liters of synthetic oil. If 344 liters of synthetic oil are used, how many liters of Petrolyn oil will be made?

OpenStudy (caozeyuan):

N=5/9, and S=4/9, ok?

OpenStudy (ashleyy_smith):

ok

OpenStudy (caozeyuan):

In general, S=4P/9, you know S, oyu can solve for P

OpenStudy (ashleyy_smith):

where did you get the 9 from?

OpenStudy (caozeyuan):

4+5=9, cuz in 9 liter of P, there is 4L of S and 5L of N according to your question

OpenStudy (ashleyy_smith):

So it's 9? @caozeyuan

OpenStudy (raffle_snaffle):

yes, he just gave you the answer.

OpenStudy (ashleyy_smith):

774?